OpenStudy (phi):

How to simplify this \[ y= \frac{\log_e\left( \frac{x}{m} -sa \right)}{r^2} \]

OpenStudy (phi):

I did this: \[ r^2 y = \log_e\left(\frac{x}{m}-sa\right) \] \[ e^{r^2y}= \left(\frac{x}{m}-sa\right) \] \[ me^{rry}= x-mas \]

OpenStudy (baru):

Haha :D nice one

Parth (parthkohli):

ahahahahaha! Merry Christmas :)

zepdrix (zepdrix):

Hehe, I remember seeing this one last year :)) lots of fun

OpenStudy (phi):

Yes, I did not invent this. But it is cute.

OpenStudy (baru):

:D merry Christmas!!! did not think i'd get to see the human side of phi X'D

OpenStudy (irishboy123):

now i can't wait for April Fool's Day....

hartnn (hartnn):

Saw this on facebook, logged in here to post it, late by an hour! :P

OpenStudy (ikram002p):

:O